### v1

# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

Abstract
Motivated by a recent method for approximate solution of Fredholm equations of the first kind, we develop a corresponding method for a class of Fredholm equations of the second kind. In particular, we consider the class of equations for which the solution is a probability measure. The approach centres around specifying a functional whose gradient flow admits a minimizer corresponding to a regularized version of the solution of the underlying equation and using a mean-field particle system to approximately simulate that flow. Theoretical support for the method is presented, along with some illustrative numerical results.
1 Introduction
Fredholm integral equations of the second kind are defined by
$\pi(x)=\varphi(x)+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y$ (1)
for any $x\in\mathbb{R}^{d}$ and some $\lambda\in\mathbb{R}$, with $\pi$ an unknown function on $\mathbb{R}^{d}$, $\varphi$ an observed forcing on $\mathbb{R}^{d}$ and $\mathrm{k}:\ \mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ a positive integral kernel—where $\pi$ and $k$ belong to suitably regular classes. We focus on the case in which $\pi$ is a positive function with known finite integral, i.e. $\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x<\infty$, so that we can renormalize the left hand side of (1) and consider the normalized version of $\pi$ and define the new forcing as $\varphi(x)/\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x$ (which we denote by $\pi,\varphi$ too with a slight abuse of notation). These integral equations find applications in numerous contexts including econometrics [20, 12], modelling light transport for ray-tracing and related applications [52, 53] and reinforcement learning [17]. In the homogeneous case (i.e. when $\varphi\equiv 0$), solving (1) corresponds to finding the eigenfunction $\pi$ of $\mathrm{k}$ for a given eigenvalue $\lambda$; the eigen–decomposition of positive kernels $\mathrm{k}$ is normally referred to as Karhunen–Loève decomposition (see [18] for a recent review) and has applications to Gaussian process regression [55], spatial statistics [15, 49] and Markov chain Monte Carlo (MCMC) in infinite dimensional spaces [1]. When $\mathrm{k}$ is the density of a Markov kernel (i.e. $\int\mathrm{k}(x,y)\mathrm{d}y=1$) and $\lambda=1$, solving (1) is equivalent to finding an invariant measure of $\mathrm{k}$. This is useful when dealing with approximate MCMC kernels [2, 44] or tractable approximations to transition densities [7].
Henceforth we will use the same symbol to denote measures and their Lebesgue densities whenever they possess them—particularly $\pi$ and $\mathrm{k}$ such that, e.g., $\mathrm{d}\pi(y)=\pi(y)\mathrm{d}y$. To solve (1) we consider a minimization problem with target functional
$\mathcal{F}_{\alpha}(\pi)=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right),$ (2)
where $\textup{KL}\left(\mu\middle|\nu\right)=\int_{\mathbb{R}^{d}}\log((\mathrm{d}\mu/\mathrm{d}\nu)(y))\mathrm{d}\mu(y)$ denotes the Kullback–Leibler divergence between $\mu$ and $\nu$, for a given regularization parameter $\alpha>0$ and some reference measure $\pi_{0}$. Henceforth, we implicitly work under the assumption that the $\pi$ of (1) is the Lebesgue density of a probability distribution over a Euclidean space. Standard approaches normally rely upon discretization of $\pi$ and $\mathrm{k}$ through basis functions or using numerical integration (see, e.g., [5] for a survey of classical methods). These methods are particularly effective in low-dimensional settings, but they become challenging when $d$ is not small and in the case in which the solution $\pi$ is defined over an unbounded interval. In this paper we focus on the latter issue, and develop a method to solve Fredholm integral equations of the second kind on unbounded domains.
Approaches based upon loss minimization have mostly focused on maximum entropy solutions [43, 31, 34] and $\mathbb{L}^{2}$ loss [29]. However, the use of the Kullback–Leibler divergence is common in the literature on Fredholm integral equations of the first kind [28, 48, 13, 32] and in that context a functional closely related to (2) has been recently shown to have properties similar to those of Tikhonov regularization [16].
We extend the approach of [16] to equations of the second kind and derive a particle approximation to a Stochastic Differential Equation (SDE) of McKean–Vlasov type (an MKVSDE) associated with the Wasserstein gradient flow minimizing (2). This approach is particularly well-suited to solving integral equations on unbounded domains since is not based on a fixed discretization over a finite support, but on an adaptive stochastic discretization with diffusive behaviour.
The main difficulty which occurs when adapting the approach of [16] to the context of Fredholm equations of the second kind is a consequence of the presence of the unknown distribution, $\pi$, on both sides of the first KL divergence in (2): this results in an MKVSDE whose drift depends upon nested expectations w.r.t. $\pi$. As a consequence, the algorithm developed herein has higher cost w.r.t. the number of particles in our particle system than that of [16] and the fundamental difference in the drift of the MKVSDE means that novel results are required to justify our approach. We derive explicit bounds on the error of the proposed numerical algorithm which can be used to set the number of particles in the system and the time discretization step.
Finally, we demonstrate empirically that our method provides accurate solutions to integral equations of the form (1) on unbounded domains and can be applied more generally than other methods (e.g. those based on the von Neumann decomposition; see our first numerical experiment). We stress that the presence of the regularization term $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ in (2) is crucial for obtaining significant improvements in the case in which the integral operator induced by the kernel $\mathrm{k}$ is not invertible or poorly conditioned—as one might expect, regularization is critical in order to obtain good solutions to this ill-posed inverse problem.
The remainder of the manuscript is organized as follows. In Section 2 we study the minimization problem in (2) and give conditions under which it admits a unique minimizer. We derive the Wasserstein gradient flow and the corresponding McKean–Vlasov SDE in Section 3. Section 4 introduces an interacting particle system approximating the MKVSDE and a numerical scheme which approximates the continuous time dynamics. We also provide explicit error bounds on the approximation. Finally, in Section 5, we test our method on several examples and compare with other methods commonly used in the literature. Julia code to reproduce all examples is available online at https://github.com/FrancescaCrucinio/FE2kind_WGF.
1.1 Notation
[…]
$\mathbf{W}_{p}(\mu,\nu)=\left(\inf_{\gamma\in\mathbf{T}(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left\|x-y\right\|^{p}\mathrm{d}\gamma(x,y)\right)^{1/p}$ (3)
where $\mathbf{T}(\mu,\nu)=\{\gamma\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d}):\gamma(\mathsf{A}\times\mathbb{R}^{d})=\mu(\mathsf{A}),\ \gamma(\mathbb{R}^{d}\times\mathsf{A})=\nu(\mathsf{A})\ \forall\mathsf{A}\in\mathcal{B}(\mathbb{R}^{d})\}$ denotes the set of all transport plans between $\mu$ and $\nu$. In the following, we metrize $\mathcal{P}_{p}(\mathbb{R}^{d})$ with $\mathbf{W}_{p}$. For all $\nu\in\mathcal{P}^{ac}(\mathbb{R}^{d})$ we denote by $\mathrm{H}(\nu):=-\int_{\mathbb{R}^{d}}\log(\nu(y))\mathrm{d}\nu(y)$ the differential entropy of $\nu$, where with a slight abuse of notation we denote by $\nu$ its density w.r.t. the Lebesgue measure.
2 The Minimization Problem
We now introduce the minimization problem for the functional (2) and give conditions under which it is well-posed. We then highlight connections with maximum entropy methods, a popular technique to solve Fredholm integral equations in low-dimensional settings.
2.1 Stability of $\mathcal{F}_{\alpha}$
We study the properties of the functional
$\mathcal{F}_{\alpha}(\pi)=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right),$
for $\alpha\geq 0$, and establish that it admits a unique minimizer.
In addition, we consider the following assumption on the kernel $\mathrm{k}$ which guarantees finiteness of the corresponding term in $\mathcal{F}_{\alpha}$.
Assumption 1.
The density of the kernel $\mathrm{k}$ and the forcing $\varphi$ are such that $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$, $\varphi\in\mathrm{C}^{\infty}(\mathbb{R}^{d},[0,+\infty))$ and there exists $\mathtt{M}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ we have $\mathrm{k}(x,y)+\left\|\nabla\mathrm{k}(x,y)\right\|+\left\|\nabla^{2}\mathrm{k}(x,y)\right\|\leq\mathtt{M}$ and $\varphi(x)+\left\|\nabla\varphi(x)\right\|+\left\|\nabla^{2}\varphi(x)\right\|\leq\mathtt{M}$.
Assumption 1 restricts the class of kernels $\mathrm{k}$ to smooth kernels with bounded derivatives up to the second order. This might seem restrictive, however, smooth kernels are particularly challenging while discountinuous or degenerate kernels normally allow for specific solution methodologies (see, e.g., [38, Chapter 11]). In this section, we will require only that $\mathrm{k}\in\mathrm{C}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$. The control of higher order derivatives is necessary to establish convergence result for the interacting particle system (both in continuous and discrete time) and will be exploited later.
To address stability issues we also consider the following regularized functional: for any $\alpha,\eta\geq 0$ and $\pi\in\mathcal{P}(\mathbb{R}^{d})$
$\mathcal{F}_{\alpha}^{\eta}(\pi)$ $=\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right),$ (4)
where $\eta\geq 0$ is a hyperparameter and $\mathcal{F}^{\eta}(\pi):=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y+\eta\right)$. In the case $\eta=0$, $\mathcal{F}_{\alpha}^{\eta}(\pi)$ coincides with $\mathcal{F}_{\alpha}(\pi)$. However, establishing the uniqueness of the minimizer of $\mathcal{F}_{\alpha}$ is difficult. In what follows we study the function $\mathcal{F}_{\alpha}^{\eta}$ for $\alpha,\eta>0$ and show that $\mathcal{F}^{\eta}_{\alpha}$, restricted to $\mathcal{P}^{ac}(\mathbb{R}^{d})$, is coercive in this case. Hence, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$.
Proposition 1.
Under Assumption 1, we have the following
  1. (a)
    For any $\eta\geq 0$, $\mathcal{F}^{\eta}:\mathcal{P}^{ac}(\mathbb{R}^{d})\to\mathbb{R}$ is lower bounded and convex.
  2. (b)
    For any $\alpha,\eta>0$, $\mathcal{F}_{\alpha}^{\eta}:\mathcal{P}^{ac}(\mathbb{R}^{d})\to\mathbb{R}$ is proper, strictly convex, coercive and lower semi-continuous. In particular, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}\in\mathcal{P}^{ac}(\mathbb{R}^{d})$.
Proof.
See Appendix A. ∎
The proof of Proposition 1 shows that the condition $\alpha>0$ is sufficient to guarantee uniqueness of the minimizer, in fact we rely on the coercivity of $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ to show that $\mathcal{F}_{\alpha}^{\eta}$ is coercive. Thus, for the remainder of this work we only consider the case $\alpha,\eta>0$.
2.2 Maximum Entropy Methods
Minimizing the functional $\mathcal{F}_{\alpha}$ is related to finding the maximum entropy solution to (1) [33]. In fact, $\mathcal{F}_{\alpha}$ can be seen as the Lagrangian associated with the following primal problem
$\arg\min\left\{\textup{KL}\left(\pi\middle|\pi_{0}\right):\pi\in\mathcal{P}(\mathbb{R}^{d}),\ \textup{KL}\left(\pi\middle|\varphi+\lambda\int\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)=0\right\}.$ (5)
To highlight the connection with maximum entropy methods, we observe that if we replace the KL penalty with an entropic penalty, we obtain a functional
~$\tilde{\mathcal{F}}_{\alpha}(\pi)=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)-\alpha\mathrm{H}(\pi),$ (6)
corresponding to the Lagrangian associated with the following primal problem
$\arg\max\left\{\mathrm{H}(\pi):\pi\in\mathcal{P}_{\mathrm{H}}(\mathbb{R}^{d}),\ \textup{KL}\left(\pi\middle|\varphi+\lambda\int\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)=0\right\},$ (7)
where $\mathcal{P}_{\mathrm{H}}(\mathbb{R}^{d})$ is the set of probability distributions with finite entropy. However, the functional ~$\tilde{\mathcal{F}}_{\alpha}$ is not lower bounded and the corresponding minimization problem is not well-defined and therefore is not considered here.
Maximum entropy solutions to Fredholm integral equations were originally proposed in [43] under moment constraints obtained by selecting different families of basis functions [34, 31]. In the one dimensional case, the maximum entropy solution can be written analytically (see, e.g., [31, Proposition 3.1]), and these approaches are particularly effective. However, in higher dimensions maximum entropy methods generally require discretization of the support of the solution $\pi$.
3 A Wasserstein Gradient Flow for $\mathcal{F}_{\alpha}^{\eta}$ and an associated McKean–Vlasov SDE
Having established that the regularized functional (4) admits a unique minimizer, we exploit the connection between minimization of functionals in the space of probability measures and partial differential equations (PDEs) identified in [35, 46] to obtain an SDE whose invariant measure can be related to the minimizer of $\mathcal{F}_{\alpha}^{\eta}$.
We start by deriving the Wasserstein gradient flow of $\mathcal{F}_{\alpha}^{\eta}$ and, after observing that the corresponding PDE is a Fokker–Plank equation, use standard stochastic calculus to obtain a McKean–Vlasov SDE (MKVSDE) whose law satisfies the Fokker–Plank PDE.
3.1 Wasserstein Gradient Flow
To derive the Wasserstein gradient flow PDE for $\mathcal{F}_{\alpha}^{\eta}$ we first derive the Wasserstein subdifferential of the functional [4, Definition 10.1.1], to guarantee the existence of this operator we further assume that $\pi_{0}$ satisfies the following assumption.
Assumption 2.
The following hold:
  1. (a)
    $\pi_{0}$ admits a density w.r.t. the Lebesgue measure, $\mathrm{d}\pi_{0}(x)=\pi_{0}(x)\mathrm{d}x$, with $\pi_{0}(x)\propto\exp\left[-U(x)\right]$, where, $U:\ \mathbb{R}^{d}\to\mathbb{R}$ belongs to $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R})$.
  2. (b)
    There exists $\mathtt{L}\geq 0$ such that $\left\|\nabla U(x_{1})-\nabla U(x_{2})\right\|\leq\mathtt{L}\left\|x_{1}-x_{2}\right\|$, for any $x_{1},x_{2}\in\mathbb{R}^{d}$.
  3. (c)
    There exist $\mathtt{m},\mathtt{c}>0$ such that for any $x_{1},x_{2}\in\mathbb{R}^{d}$,
    $\langle\nabla U(x_{1})-\nabla U(x_{2}),x_{1}-x_{2}\rangle\geq\mathtt{m}\left\|x_{1}-x_{2}\right\|^{2}-\mathtt{c}.$
In Appendix B we show that under Assumption 1 and 2(a) the subdifferential of $\mathcal{F}_{\alpha}^{\eta}$ is given by
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ =$=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right.$
$\left.\qquad\qquad+\nabla\log\pi(x)+\alpha\nabla\log(\mathrm{d}\pi/\mathrm{d}\pi_{0})(x)\right\},$
where we defined $\pi[\mathrm{k}(z,\cdot)]:=\int_{\mathbb{R}^{d}}\mathrm{k}(z,y)\pi(y)\mathrm{d}y$.
The Wasserstein gradient flow [4, Definition 11.1.1] associated with $\mathcal{F}_{\alpha}^{\eta}$ is a family of probability measures $(\pi_{t})_{t\geq 0}$ satisfying (in a weak sense) $\partial_{t}\pi_{t}=\nabla\cdot(\pi_{t}\ \partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi))$, which in our case corresponds to:
$-\nabla\cdot\left(\pi_{t}\left\{\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi_{t}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi_{t}\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi_{t}\left(z\right)-\alpha\nabla U(x)\right\}\right)$ (8)
$\qquad\qquad+(1+\alpha)\triangle\pi_{t}.$
For strongly geodesically convex (i.e. convex along geodesics) functionals the Wasserstein gradient flow converges geometrically towards the unique minimizer. In our setting $\mathcal{F}_{\alpha}^{\eta}$ is not geodesically convex but only convex; in the following section we will show that $\pi_{t}$ in (8) converges to a limiting measure, $\pi_{\alpha,\eta}^{\star}$.
3.2 McKean–Vlasov SDE
Since (8) is a Fokker–Plank equation, we informally derive the corresponding SDE, see, e.g. [6] for the precise derivation,
$\mathrm{d}\bm{X}_{t}=\left\{\int b^{\eta}(\bm{X}_{t},z,\pi_{t})\mathrm{d}\pi_{t}(z)-\alpha\nabla U(\bm{X}_{t})\right\}\mathrm{d}t+\sqrt{2(1+\alpha)}\mathrm{d}\bm{B}_{t}$ (9)
where $(\bm{B}_{t})_{t\geq 0}$ is a $d$-dimensional Brownian motion, $\pi_{t}$ is the law of $(\bm{X}_{t})_{t\geq 0}$ and for any $\nu\in\mathcal{P}(\mathbb{R}^{d})$ and $(x,z)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$
$b^{\eta}(x,z,\nu)=\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\nu\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}.$ (10)
The presence of $\eta>0$ in the denominator guarantees that $b^{\eta}$ is always well defined.
The SDE (9) belongs to the class of McKean–Vlasov SDEs since the drift coefficient depends upon not only $\bm{X}_{t}$ but also its law, $\pi_{t}$. Under Assumption 1 and 2, if $\eta>0$ the drift coefficient of (9) is Lipschitz continuous and we can establish existence and uniqueness for the MKVSDE.
Proposition 2.
Under Assumptions 1, 2(a) and 2(b), for any $\alpha,\eta>0$ there exists a unique strong solution to (9) for any initial condition $\bm{X}_{0}$ such that $\mathcal{L}(\bm{X}_{0})\in\mathcal{P}^{ac}_{1}(\mathbb{R}^{d})$.
Proof.
See Appendix C.1. ∎
The previous proposition is limited to the case where $\eta>0$. Indeed, if $\eta=0$ the drift is not Lipschitz continuous and the SDE (9) might be unstable, with solutions existing up to a (possibly small) explosion time. Using recent results from [30] we can establish that $\pi_{t}$ converges to the unique minimizer $\pi_{\alpha,\eta}^{\star}$ of $\mathcal{F}_{\alpha}^{\eta}$ when $t\to+\infty$. The presence of an interaction term $b^{\eta}$ which is not small w.r.t. $\alpha U$ prevents us from using more standard approaches (e.g. [11, 23, 41, 9]) and thus from obtaining quantitative convergence rates.
Proposition 3.
Under Assumptions 1 and 2, for any $\alpha,\eta>0$, we have
$\underset{t\to+\infty}{\lim}\mathbf{W}_{2}(\pi_{t},\pi_{\alpha,\eta}^{\star})=0.$
Proof.
See Appendix C.2. ∎
The result above can be strengthened to show that, if $\alpha_{t}\to 0$ at the appropriate rate, then $\mathcal{F}^{\eta}(\pi_{t})$ converges towards the minimum of the unregularized functional (see [14, Theorem 4.1] and [45, Theorem 1]).
4 Particle System
As discussed in the previous section, the drift of the MKVSDE (9) cannot be computed analytically, since it depends on the law $\pi_{t}$ of $\bm{X}_{t}$ at each time step. Thus, it is not possible to directly approximate (9) with a time-discretized process.
A classical approach to circumvent this issue is to consider an interacting particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}=\{(\bm{X}_{t}^{i,N})_{t\geq 0}\}_{i=1}^{N}$ for any $N\in\mathbb{N}^{\star}$ which approximates (9) and satisfies a classical SDE [42, 10]. We introduce the particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}$ which satisfies the following SDE: for any $i\in\{1,\dots,N\}$, $\bm{X}_{0}^{i,N}\in\mathbb{R}^{d}$ and
$\mathrm{d}\bm{X}_{t}^{i,N}=\left\{\int b^{\eta}(\bm{X}_{t}^{i,N},z,\pi_{t}^{N})\mathrm{d}\pi_{t}^{N}\left(z\right)-\alpha\nabla U(\bm{X}_{t}^{i,N})\right\}\mathrm{d}t+\sqrt{2(1+\alpha)}\mathrm{d}\bm{B}_{t}^{i},$ (11)
where $\{(\bm{B}_{t}^{i})_{t\geq 0}\}_{i\in\mathbb{N}}$ is a family of independent Brownian motions and $\pi_{t}^{N}$ is the empirical measure associated with $(\bm{X}_{t}^{1:N})_{t\geq 0}$, $\pi_{t}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{X}_{t}^{i,N}}$ for every $t\geq 0$.
In the following proposition we establish a classical propagation of chaos result, showing that there exists a unique strong solution to (11) and that this solution approximates (9) for any finite time horizon.
Proposition 4.
Under Assumptions 1, 2(a) and 2(b) for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exists a unique strong solution to (11) for any initial condition $\bm{X}_{0}^{1:N}$ such that $\mathcal{L}(\bm{X}_{0}^{1:N})\in\mathcal{P}_{1}^{ac}((\mathbb{R}^{d})^{N})$ and $\{\bm{X}_{0}^{i,N}\}_{i=1}^{N}$ is exchangeable. In addition, for any $T\geq 0$ there exists $c_{1}(T)\geq 0$ such that for any $N\in\mathbb{N}^{\star}$ and $\ell\in\{1,\dots,N\}$ if $\bm{X}_{0}^{\ell,N}=\bm{X}_{0}$ almost surely, then:
$\mathbb{E}\left[\sup_{t\in[0,T]}\left\|\bm{X}_{t}-\bm{X}_{t}^{\ell,N}\right\|^{2}\right]\leq\frac{c_{1}(T)}{\sqrt{N}}$
Proof.
See Appendix D.1. ∎
To obtain stronger convergence results for our particle system we use the dissipativity condition in Assumption 2(c). This dissipativity condition allows us to obtain the exponential ergodicity of the particle system using standard tools in the literature on convergence of SDEs. The bounds that we obtain, however, are not uniform w.r.t to the number of particles and, in particular, they depend on a constant $C_{N}\to\infty$ and a rate $\rho_{N}\to 1$ as $N\to+\infty$.
Proposition 5.
Under Assumptions 1 and 2, for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exist $C_{N}\geq 0$ and $\rho_{N}\in[0,1)$ such that for any $x_{1}^{1:N},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
$\mathbf{W}_{1}(\pi_{t}^{N}(x_{1}^{1:N}),\pi_{t}^{N}(x_{2}^{1:N}))\leq C_{N}\rho_{N}^{t}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|,$
where for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$, $\pi_{t}^{N}(x^{1:N})$ is the distribution of $(\bm{X}_{t}^{1:N})_{t\geq 0}$ with initial condition $x^{1:N}$. In particular, (11) admits a unique invariant probability measure denoted $\pi^{N}\in\mathcal{P}_{p}((\mathbb{R}^{d})^{N})$ and for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
$\mathbf{W}_{1}(\pi_{t}^{N}(x^{1:N}),\pi^{N})\leq C_{N}\rho_{N}^{t}\left(\left\|x^{1:N}\right\|+\int_{\mathbb{R}^{d}}\left\|\tilde{x}\right\|\mathrm{d}\pi^{N}(\tilde{x})\right).$
Proof.
See Appendix D.2. ∎
Proposition 4 and 5 both show that the interacting particle system (11) is well behaved: Proposition 4 shows that for any finite time horizon $T$ (11) well approximates (9), whereas, Proposition 5 guarantees that (11) has a unique invariant measure for each finite $N$. The next result guarantees that the invariant measure $\pi^{N}$ of (11) converges to $\pi_{\alpha,\eta}^{\star}$.
Proposition 6.
Under Assumptions 1 and 2, for any $\alpha,\eta>0$,
$\lim_{N\to+\infty}\mathbf{W}_{1}(\pi^{N},\pi_{\alpha,\eta}^{\star})=0.$
In addition, $\pi_{\alpha,\eta}^{\star}$ is the unique invariant probability measure of (9).
Proof.
See Appendix D.3. ∎
4.1 Euler–Maruyama Discretization
Having obtained a particle system (11) whose invariant distribution converges to the minimizer $\pi_{\alpha,\eta}^{\star}$ of (2), we now consider the numerical approximation of (11) using an Euler–Maruyama time discretization scheme. For any $N\in\mathbb{N}$, we consider the following time discretization given by $X_{0}^{1:N}\in(\mathbb{R}^{d})^{N}$ and for any $n\in\mathbb{N}$ and $k\in\{1,\dots,N\}$
$X_{n+1}^{k,N}=X_{n}^{k,N}+\gamma b(X_{n}^{k,N},\pi^{N}_{n})+\sqrt{2\gamma(\alpha+1)}Z_{n+1}^{k},$ (12)
where $\{Z_{n}^{k}\}_{k,n\in\mathbb{N}}$ is a family of independent standard Gaussian random variables, $\gamma>0$ is a stepsize, $b$ is given by
$b(x,\pi):=\int_{\mathbb{R}^{d}}b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x),$
and for any $n\in\mathbb{N}$, we have that $\pi^{N}_{n}=(1/N)\sum_{k=1}^{N}\delta_{X_{n}^{k,N}}$.
Since under our assumptions the drift $b$ is Lipschitz continuous in both arguments (see Appendix C.1 for a proof), [10] guarantees that the Euler–Maruyama scheme (12) has strong order of convergence $1/2$: for any $N\in\mathbb{N}^{\star}$, $\ell\in\{1,\dots,N\}$, $\gamma>0$ and $T\geq 0$, there exists $c_{2}(T)\geq 0$, independent of $\gamma$, such that
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\ell,N}-X_{n}^{\ell,N}\right\|\right]\leq c_{2}(T)\sqrt{\gamma}$ (13)
where $n_{T}=\lfloor T/\gamma\rfloor$.
For small values of $\gamma>0$ and large values of $n\gamma$ and $N\in\mathbb{N}$ we get that $\pi_{n}^{N}=(1/N)\sum_{k=1}^{N}\delta_{X_{n}^{k,N}}$ is an approximation of $\pi_{\alpha,\eta}^{\star}$. To obtain a smooth approximation, we can plug in this particle approximation into (1) and define $\hat{\pi}_{n}^{N}:\ \mathbb{R}^{d}\to\mathbb{R}$ such that for any $x\in\mathbb{R}^{d}$
$\hat{\pi}_{n}^{N}(x)=\varphi(x)+\frac{\lambda}{N}\sum_{k=1}^{N}\mathrm{k}(x,X_{n}^{k,N}),$ (14)
as suggested in [21] as a method for obtaining smooth approximations of the solution of a Fredholm equation from a particle approximation. Our final algorithm is summarized in Algorithm 1. For simplicity we assume initialization is carried out by simple random sampling from some initial distribution but other possibilities, including stratified sampling and using quasi Monte Carlo pointsets might further improve performance.
Algorithm 1 FE2kind-WGF: a method for solving Fredholm integral equations of the second kind with Wasserstein gradient flows.
  Require: $N,n_{T}\in\mathbb{N}$, $\alpha,\eta,\gamma>0$, $\mu,\pi_{0},\pi_{\mathrm{init}}\in\mathcal{P}^{ac}(\mathbb{R}^{d})$.
  Draw $\{X_{0}^{k,N}\}_{k=1}^{N}$ from $\pi_{\mathrm{init}}^{\otimes N}$
  for $n=1:n_{T}$ do
     for $k=1:N$ do
        Update $X_{n}^{k,N}=X_{n-1}^{k,N}+\gamma b(X_{n-1}^{k,N},\pi^{N}_{n-1})+\sqrt{2\gamma(\alpha+1)}Z^{k}_{n}$ as in (12)
     end for
  end for
  return  $\hat{\pi}_{n}^{N}(x)$ as in (14)
4.2 Implementation Guidelines
Efficient Computation
Notice that the drift of (9) involves nested expectations w.r.t. $\pi_{t}$. In particular, we can decompose the drift into
$b(x,\pi)$ $=\int b^{\eta}(x,z,\pi_{t})\mathrm{d}\pi_{t}(z)-\alpha\nabla U(x)$
$=\int\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\pi(z)+\frac{\lambda\pi[\nabla_{1}\mathrm{k}(x,\cdot)]+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}-\alpha\nabla U(x).$
Naïve implementation of the evaluation of $b$ might lead to a $\mathcal{O}(N^{3})$ cost. However, precomputing $\pi^{N}_{n}\left[\mathrm{k}(X_{n}^{j,N},\cdot)\right]$ and storing its value for each $j=1,\dots,N$ allows computation at an $\mathcal{O}(N^{2})$ cost.
Choice of $\pi_{0}$
The reference measure $\pi_{0}$ is the mechanism by which the choice of regularization is specified and can be chosen to impose particular properties (e.g. degree of smoothness, localized support) on $\pi$ or to favour solutions with such properties. While the presence of $\pi_{0}$ might seem arbitrary, (1) is an inverse problem, and thus regularization is required to obtain stable results. This manifests itself in the fact that $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer only when $\alpha>0$ (Proposition 1).
The choice of an improper reference measure $\pi_{0}\propto C$ with $C>0$ might seem a natural pragmatic choice. However, this results in a MKVSDE whose drift does not depend on $\pi_{0}$
$\mathrm{d}\bm{X}_{t}=\left\{\int b^{\eta}(\bm{X}_{t},z,\pi_{t})\mathrm{d}\pi_{t}(z)\right\}\mathrm{d}t+\sqrt{2(1+\alpha)}\mathrm{d}\bm{B}_{t}.$ (15)
This scheme corresponds to constructing the gradient flow for the functional (6), that, as previously discussed, does not necessarily lead to a unique minimizer. In addition, our proof of convergence of $\pi_{t}$ to the unique minimizer (Proposition 3) relies on the fact that $\alpha>0$ with som $\pi_{0}$ satisfying Assumption 2.
In our first experiment we show how the presence of $\pi_{0}$ stabilizes the minimizer $\pi_{\alpha,\eta}^{\star}$. See also [16, Appendix D.1] for a discussion on how $\pi_{0}$ might influence the results in a similar context.
Tuning and algorithmic set up
The value of $\alpha$ controls the amount of regularization introduced by the cross-entropy penalty, and, as shown by Proposition 1, when $\alpha>0$ the functional $\mathcal{F}_{\alpha}^{\eta}$ is coercive and thus admits a unique minimizer. The value of $\alpha$ could be fixed a priori, or, to guarantee a good trade-off between accuracy and regularization, it could be selected by cross-validation as suggested in [16, 54, 3] for Fredholm integral equations of the first kind.
The parameter $\eta$ was introduced in (4) to deal with the possible instability of the functional $\mathcal{F}_{\alpha}^{\eta}$; we did not find performances to be significantly influenced by this parameter as long as its value is sufficiently small. In practice, in the experiments in Section 5 we set $\eta\equiv 0$ and observe that the resulting algorithm is stable. This was also empirically observed in related contexts by [16] where a tamed Euler discretization is employed, and by [40] where a standard Euler discretization is employed.
The values of the number of particles $N$ and the time discretization step $\gamma$ control the quality of the numerical approximation of (9). Combining the result in Proposition 4 with (13) we obtain the following global error estimate: for any $T\geq 0$ we have
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\star}-X_{n}^{\ell,N}\right\|\right]\leq\frac{c_{1}(T)}{\sqrt{N}}+c_{2}(T)\sqrt{\gamma}$ (16)
Choosing $N$ amounts to the classical task of selecting an appropriate sample size for Monte Carlo approximations, while the choice of $\gamma$ corresponds to the specification of a timescale on which to discretize a continuous time (stochastic) process; hence, one can exploit the vast literature on Monte Carlo methods and discrete time approximations of SDEs to select these values [37].
Given a fixed computational budget $B$, and observing that the cost of running Algorithm 1 is $\mathcal{O}(N^{2}/\gamma)$, we can minimize the r.h.s. of (16) subject to $B=N^{2}/\gamma$ and obtain, via a simple Lagrange multiplier argument, that the optimal values of $N,\gamma$ are given by $N=B^{1/3}[c_{1}(T)/(2c_{2}(T))]^{2/3}$ and $\gamma=N^{2}/B=(c_{1}(T)/2c_{2}(T))^{4/3}B^{-1/3}$. This suggests that, optimally, $N$ and $1/\gamma$ both scale with the cube root of the computational effort available—suggesting setting $\gamma=O(N^{-1})$, which results in an error decaying at the usual Monte Carlo rate of $O(N^{-1/2})$ with an overall cost scaling as $O(N^{3})$.
It is straightforward to choose the number of time steps $n_{T}$ adaptively by approximating the value of $\mathcal{F}_{\alpha}^{\eta}$ through numerical integration by replacing the r.h.s. of (1) with (14) and using standard kernel density estimation [50] for the l.h.s.. Once the value of $\mathcal{F}_{\alpha}^{\eta}$ stops decreasing, a minimizer has apparently been reached and the iteration can be stopped.
5 Experiments
5.1 Simple Gaussian Example
We start by considering a simple integral equation in which the solution $\pi$ can be computed analytically. Consider (1) with $\varphi(x)=(1-\lambda)\mathcal{N}\left(x;0,1\right)$ and $\mathrm{k}(x,y)=\mathcal{N}\left(y;xe^{-\beta},(1-e^{-2\beta})\right)$ for some $10<\lambda<$ 1, $\beta>0$, with $\mathcal{N}(x;\mu,\sigma^{2})$ denoting the density of a normal distribution of mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$. Since $\max_{x}\int_{\mathbb{R}^{d}}\lambda k(x,y)\mathrm{d}y=\lambda<1$, [38, Corollary 2.16] guarantees that (1) admits a unique solution given by $\pi(x)=\mathcal{N}\left(x;0,1\right)$ for all $10<\lambda<$ 1. In our experiments we set $\beta=0.5$.
5.1.1 Effect of Reference Measure
As discussed in Section 4.2, the reference measure $\pi_{0}$ can be used to impose certain properties on the regularized solution $\pi_{\alpha,\eta}^{\star}$. To check the importance of using a reference measure, we compare the approximate solutions obtained with an improper reference measure (i.e. one corresponding to the entropic penalty in (6)) and several choices of $\pi_{0}$: the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We set $N=100$, $\gamma=10^{-2}$ and iterate for $n_{T}=200$, which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). We consider $\alpha\in[0,1]$.
We check the accuracy of the reconstructions through their mean, variance and integrated square error
$\operatorname{ISE}(\hat{\pi})=\int_{\mathbb{R}}\{\pi(x)-\hat{\pi}(x)\}^{2}\mathrm{d}x,$ (17)
with $\hat{\pi}$ an estimator of $\pi$ (the integral is approximated by numerical integration).
[…]
Figure 1: Effect of reference measure on reconstruction accuracy as $\alpha$ increases. We compare $\operatorname{ISE}$ and $\operatorname{MSE}$ of both mean and variance.
Figure 1 shows the $\operatorname{ISE}$ and the mean squared error ($\operatorname{MSE}$) for mean and variance of the reconstructions as the value of the regularizing parameter $\alpha$ increases. Large values of $\alpha$ inflate the diffusion coefficient of the MKVSDEs (9) and (15), but, in the latter, there is no $\nabla U$ term in the drift and thus the diffusion term becomes stronger than the drift term for large $\alpha$s, causing the reconstructed solutions to have larger variance than the solution. A reference measure more diffuse than the target has a similar effect.
It may seem superficially surprising that a concentrated measure seems to give better results than using $\pi_{0}=\pi$ if one looks at the mean alone, but this simply reflects the fact that a concentrated measure forces the particles to be close to its mean (in this case 0) and therefore compensates for the effect of the diffusion term in the SDE (9). However, a concentrated measure performs worse in terms of variance, especially for large $\alpha$.
5.1.2 Comparison with Reversible jump MCMC
The previous experiments show that the presence of the reference measure significantly improves the results provided by FE2kind-WGF. We now compare our approach with the trans-dimensional MCMC algorithm proposed in [21] which provides a Monte Carlo approximation of the von Neumann series representation of the solution $\pi$ (see, e.g., [38, Section 2.4]). We implement their reversible jump MCMC (RJ-MCMC) with the following set up: we set the death and birth probabilities to 1/3 and use a random walk proposal with variance $\sigma^{2}=0.1^{2}$ for the update move and sample from $\pi$ for the birth move.
For FE2kind-WGF we set $\alpha=0.01$, $N=100$ and $\gamma=10^{-2}$ and iterate for $n_{T}=200$ time steps. We compare three reference measures, the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We do not consider the case in which no reference measure is given further as the experiments in the previous section showed that this approach does not perform well. The initial distribution of the particles is a Gaussian with mean 0 and small variance ($\sigma^{2}=0.1^{2}$).
We compare the behaviour of the two approaches as $\lambda$ varies in $(0,1)$. We set the number of sample paths drawn using RJ-MCMC to $M=2\cdot 10^{4}$ to match the average cost of FE2kind-WGF (roughly 0.25 seconds for both algorithms). A smooth reconstruction is obtained by using (14) for FE2kind-WGF and using an equivalent strategy for the RJ-MCMC estimator (see [21, Eq. (44)]).
Figure 2 shows the average $\operatorname{ISE}$ and the mean squared error of the estimated variance over 100 replicates for $\lambda\in(0,1)$. As $\lambda\to 1$ recovering the solution becomes more challenging, as the von Neumann series becomes unstable. However, the results provided by FE2kind-WGF remain stable as $\lambda$ increases for all choices of $\pi_{0}$, and in fact, can be applied even when $\lambda=1$ (i.e. no forcing term; see Section 5.3), while RJ-MCMC cannot. Indeed, FE2kind-WGF provides reconstructions as least as good as those of the RJ-MCMC approach for all values of $\lambda$ and has considerably better behaviour as $\lambda$ increases. This is not surprising as the larger the value of $\lambda$, the more terms of the von Neumann expansion that make significant contributions and so the larger the space that the RJ-MCMC algorithm is required to explore in order to give a good approximation overall whereas the FE2kind-WGF approach does not depend upon such an expansion.
[…]
Figure 2: Accuracy of solutions as $\lambda$ increases on a toy Gaussian model. We compare FE2kind-WGF and RJ-MCMC with similar cost through the $\operatorname{ISE}$ and the $\operatorname{MSE}$ of the estimated variance.
5.2 Karhunen–Loève Expansions
The Karhunen–Loève expansion of a stochastic process with covariance function $\mathrm{k}(x,y)$ is given by the pairs of eigenvalues and eigenfunctions $(\lambda,\pi)$ of $\mathrm{k}$. For each eigenvalue $\lambda$, the corresponding eigenfunction can be found by solving (1) with $\varphi\equiv 0$. We consider two examples commonly used in the Gaussian process literature [56, Section 4.2]: the squared exponential kernel, $\mathrm{k}(x,y)=\exp(-(y-x)^{2})$, which satisfies our assumptions, and the exponential kernel, $\mathrm{k}(x,y)=\exp(-|y-x|)$, which is not differentiable for $y=x$ and therefore does not satisfy Assumption 1.
The Karhunen–Loève expansion of the exponential kernel is known analytically [27, Section 2.3.3]; the largest eigenvalue is $\lambda=2/(1+\omega^{2})$, where $\omega$ is the largest positive root of $f(\omega)=1-\omega\tan(\omega)$, with corresponding eigenfunction $\pi(x)=\cos(\lambda x)/\sqrt{1+\sin(2\lambda)/(2\lambda)}$ over $[-1,1]$.
We compare the results obtained with the Nyström method, i.e. solving the eigenvalue problem associated with the matrix obtained by discretizing $\mathrm{k}$ over the interval $[-1,1]$, with the results obtained using FE2kind-WGF. We compare the results of the two methods for increasing precision, corresponding to the number of particles $N$ for FE2kind-WGF and the number of discretization intervals for the Nyström method. In particular, we consider $N$ and the number of discretization intervals between 50 and $10^{3}$; for FE2kind-WGF we set $\gamma=1/N$, $\alpha=10^{-2}$ and iterate for $n_{T}=400$. The initial distribution and the reference measure $\pi_{0}$ are both Gaussians centred at 0 and with small variance ($\sigma^{2}=0.05^{2}$).
Since the Nyström method is a deterministic algorithm we take its $\operatorname{ISE}$ as reference and investigate the gains obtained by using FE2kind-WGF, see Figure 3 first panel. FE2kind-WGF produces results up to 5 times more accurate than those of the Nyström method for large $N$, while still outperforming the latter for small $N$.
The second and third panel of Figure 3 show the eigenfunction associated with the largest eigenvalue for the exponential and squared exponential kernel obtained with $N=500$ for FE2kind-WGF and 500 discretization intervals for the Nyström method. In the case of the squared exponential kernel, the $\lambda$ used in Algorithm 1 is that returned by the Nyström method. The two methods return coherent approximations of the eigenfunction, but our approach does not require a fixed space discretization and has considerable lower error.
[…]
Figure 3: Comparison of Nyström method and FE2kind-WGF to reconstruct the first eigenfunction of common covariance kernels. Left panel: distribution of $\operatorname{ISE}$ ratios ($\operatorname{ISE}$ of Nyström method divide by $\operatorname{ISE}$ of FE2kind-WGF). Middle and right panel: approximation of the eigenfunction of the largest eigenvalue for the exponential and squared exponential kernel.
5.3 Equilibrium Distribution of Gaussian Process State Space Models
Finally, we consider an example in which the forcing term $\varphi(x)\equiv 0$ and $\lambda=1$ and solving the integral equation (1) is equivalent to finding the invariant measure of the kernel $\mathrm{k}$ (or equivalently, the eigenmeasure corresponding to eigenvalue $\lambda=1$). In this case, the Von Neumann series does not provide a non-trivial solution, and the approach of [21] cannot be applied.
[7] analyses the equilibrium distribution of Gaussian process state space models (GP-SSM) for a one-dimensional SSM with transition function $f(x)=0.01x^{3}-0.2x^{2}+0.2x$. We fit a GP to learn $f$ using $m$ training pairs, $D_{m}:=(x_{i},f(x_{i})+\epsilon_{i})_{i=1}^{m}$, where $\epsilon_{i}$ are independent Gaussians with mean 0 and standard deviation 5. We follow [7] and use $m=20$ input points $(x_{i})_{i=1}^{m}$ uniformly distributed in $[-5,5]$, and use a squared exponential covariance function $c(x,x^{\prime})=\sigma_{f}^{2}\exp\left(-\left\|x-x^{\prime}\right\|^{2}/(2\ell^{2})\right)$, with $\ell^{2}=3.59^{2}$ and $\sigma_{f}^{2}=4.21^{2}$.
The predictive distribution of the GP-SSM is given by a Fredholm integral equation (1) with $\mathrm{k}(x,y)=\mathcal{N}(x;\mu(y,D_{m}),\sigma^{2}(y,D_{m}))$, where
$\mu(y,D_{m})$ $=c(y,x_{1:m})^{T}(c(x_{1:m},x_{1:m})+\operatorname{Id})^{-1}z_{1:m}$
$\sigma^{2}(y,D_{m})$ $=c(y,y)-c(y,x_{1:m})^{T}(c(x_{1:m},x_{1:m})+\operatorname{Id})^{-1}c(y,x_{1:m}),$
with $x_{1:m}:=(x_{1},\dots,x_{m})$ and $z_{1:m}:=(f(x_{1})+\epsilon_{1},\dots,f(x_{m})+\epsilon_{m})$. We compare the results obtained with FE2kind-WGF with that given by the Nyström method described in [7]. To ensure that the solution is non-trivial, i.e. $\pi(x)\neq 0$ for some $x$, we solve the linear system given by the Nyström method using least-squares with the additional constraint that the solution should be a probability density. FE2kind-WGF automatically enforces this constraint.
For the Nyström method we use 500 nodes in $[-20,10]$. For FE2kind-WGF we use $N=200,\gamma=0.005$ and iterate for $n_{T}=100$ iterations. The reference measure is a Gaussian with mean 0 and standard deviation 1, we set $\alpha=0.001$. Figure 4 shows the predictive distributions obtained with the two methods. To assess the quality of the results obtained we sample from the predictive distributions $\pi$ obtained with the two algorithms $n=2\cdot 10^{4}$ states $X_{1},\dots,X_{n}$ and apply the transformation $Y_{i}=\mu(X_{i},D_{m})+\sigma(X_{i},D_{m})\xi$, where $\xi$ is a standard Gaussian random variable. To verify that the obtained $\hat{\pi}$ is indeed an invariant distribution we compare our approximations of $\pi$ with the histogram of the $Y_{i}$s, the fit is good for both algorithms (right panel).
[…]
Figure 4: Predictive distribution for a 1-dimensional GP-SSM. Left: distribution recovered by FE2kind-WGF and Nyström. Right: Histogram of $x_{k+1}$ obtained by sampling $x_{k}$ from the predictive distribution provided by FE2kind-WGF.
6 Discussion
In this paper we extended the approach of [16] to Fredholm integral equations of the second kind. Under similar assumptions to those in [16] on the kernel $\mathrm{k}$, we show that the regularized functional (4) admits a unique minimizer which is the limiting distribution of the McKean–Vlasov SDE (9).
We employ an interacting particle system (11) to approximate this SDE and derive ergodicity and propagation of chaos results. Combining the latter with strong convergence results of Euler–Maruyama schemes we obtain the bound (16) which provides a practical guideline for the selection of the time discretization step $\gamma$ and the number of particles $N$.
We show that our method is able to solve a wide variety of equations of the second kind and is competitive with standard methods based on deterministic discretizations. As in the case of [16], the presence of the reference measure $\pi_{0}$ proves beneficial when the problem difficulty increases (Section 5.1). In addition, FE2kind-WGF is robust to small deviations from Assumption 1, as shown in Section 5.2 where we consider a kernel $\mathrm{k}$ which is not continuously differentiable.
While we focused on one dimensional examples to benchmark our method against other algorithms proposed in the literature, we believe that FE2kind-WGF has the potential to tackle higher dimensional problems and, as shown in [16] for equations of the first kind, to significantly outperform methods based on deterministic discretizations. One area in which this could prove particularly beneficial is spatial statistics, in which multidimensional spatial Karhunen–Loève expansions are often employed [26, 15].
One advantage of the Nyström method against FE2kind-WGF is that the latter requires knowledge of the eigenvalue $\lambda$ for which we seek to find the corresponding eigenfunction. One possible way to relax this requirement is to modify the functional (6) to $\mathcal{F}_{\alpha}(\pi,\lambda)$, allowing the definition of a minimization problem over the product space $\mathcal{P}(\mathbb{R}^{d})\times\mathbb{R}^{d}$ as explored in [39] in a different setting.
Acknowledgments
AMJ acknowledges financial support from the Engineering and Physical Sciences Research Council (EPSRC; grants EP/R034710/1 and EP/T004134/1) and by United Kingdom Research and Innovation (UKRI) via grant EP/Y014650/1, as part of the ERC Synergy project OCEAN. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Data access statement: No new data was created during this research. Julia code to reproduce all examples is available online at https://github.com/FrancescaCrucinio/FE2kind_WGF.
Appendix A Proof of Proposition 1
Before proceeding to the proof of Proposition 1 we state and prove the following auxiliary result which is adapted from [25, Appendix A.1].
Lemma 1.
Let $\nu,\mu$ be $\sigma$-finite and finite measures, respectively, on $\mathbb{R}^{d}$, and denote by $\mathcal{M}^{+}(\mathbb{R}^{d})$ the set of $\sigma$-finite measures on $\mathbb{R}^{d}$. Assume that $\mu\ll\nu$. Then, the Kullback–Leibler divergence $\textup{KL}\left(\mu\middle|\nu\right)=\int\limits_{\mathbb{R}^{d}}\log((\mathrm{d}\mu/\mathrm{d}\nu)(y))\mathrm{d}\mu(y)$ is convex in both arguments.
Proof.
The KL divergence above can be written as an $f$-divergence for $f(s)=s\log s$ with the convention $00\log 0=$ 0, $\textup{KL}\left(\mu\middle|\nu\right)=\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\mu/\mathrm{d}\nu)\mathrm{d}\nu$. The convex conjugate of $f$ is, for all $u\in\mathbb{R}$,
$f^{\star}(u):=\sup_{s>0}(su-f(s))=e^{u-1},$
and satisfies $f(s)+f^{\star}(u)\geq su$ by Fenchel’s inequality.
We will now show that the KL divergence can be expressed as the solution of a dual concave problem:
$\textup{KL}\left(\mu\middle|\nu\right)=\sup_{h\in\mathrm{C}_{b}(\mathbb{R}^{d})}\left[\int\limits_{\mathbb{R}^{d}}h(x)\mathrm{d}\mu(x)-\int\limits_{\mathbb{R}^{d}}e^{h(x)-1}\mathrm{d}\nu(x)\right],$ (18)
where $\mathrm{C}_{b}(\mathbb{R}^{d})$ denotes the set of real-valued bounded continuous functions over $\mathbb{R}^{d}$. Since the supremum in (18) is taken over a convex function of $\nu,\mu$ we will consequently obtain that $\textup{KL}\left(\mu\middle|\nu\right)$ is convex in both arguments.
We now prove (18). If $\mu$ is absolutely continuous w.r.t. $\nu$ we have
$\textup{KL}\left(\mu\middle|\nu\right)-\left[\int\limits_{\mathbb{R}^{d}}h(x)\mathrm{d}\mu(x)-\int\limits_{\mathbb{R}^{d}}e^{h(x)-1}\mathrm{d}\nu(x)\right]$
$\qquad\qquad=\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\right)\mathrm{d}\nu+\int\limits_{\mathbb{R}^{d}}f^{\star}(h(x))\mathrm{d}\nu(x)-\int\limits_{\mathbb{R}^{d}}h(x)\mathrm{d}\mu(x)\geq 0,$
by Fenchel’s inequality. This gives an upper bound on the $\sup$ in (18). We will now establish that it is also bounded below by the KL divergence.
Let us define $h^{\star}:=\log\mathrm{d}\mu/\mathrm{d}\nu+1$ and observe that
$\textup{KL}\left(\mu\middle|\nu\right)=\left[\int\limits_{\mathbb{R}^{d}}h^{\star}(x)\mathrm{d}\mu(x)-\int\limits_{\mathbb{R}^{d}}e^{h^{\star}(x)-1}\mathrm{d}\nu(x)\right].$
Then, if $h_{n}=\log\mathrm{d}\mu/\mathrm{d}\nu\mathbbm{1}_{1/n\leq\log\mathrm{d}\mu/\mathrm{d}\nu\leq n}+1\in\mathrm{C}_{b}(\mathbb{R}^{d})$, the monotone and dominated convergence theorems give
$\left[\int\limits_{\mathbb{R}^{d}}h_{n}(x)\mathrm{d}\mu(x)-\int\limits_{\mathbb{R}^{d}}e^{h_{n}(x)-1}\mathrm{d}\nu(x)\right]\to\textup{KL}\left(\mu\middle|\nu\right)$
as $n\to\infty$. Hence, we obtain a lower bound on the $\sup$ in (18).
It follows that the optimal value in (18) is bounded above and below by $\textup{KL}\left(\mu\middle|\nu\right)$ giving the result. ∎
Proof of Proposition 1.
  1. (a)
    In the case $\eta=0$ we have
    $\mathcal{F}(\pi)$ $=\textup{KL}\left(\pi\middle|\varphi+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)\geq 0$
    by definition since both $\pi$ and $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y\right)\mathrm{d}x$ are probability measures. To show that $\mathcal{F}(\pi)$ is convex, consider the functional $\mathcal{M}:\pi\mapsto\left(\varphi+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\mathrm{d}\pi(y)\right)$ which is linear w.r.t $\pi$ and therefore convex. Thus, we have that $\mathcal{F}(\pi)$ can be written as $\textup{KL}\left(\pi\middle|\mathcal{M}(\pi)\right)$. In addition, [22, Lemma 1.4.3-(b)] guarantees that the KL between probability distributions is jointly convex in both arguments. It follows that, for all $t\in[0,1]$
    $\textup{KL}\left(t\pi_{1}+(1-t)\pi_{2}\middle|\mathcal{M}(t\pi_{1}+(1-t)\pi_{2})\right)$
    $\qquad=\textup{KL}\left(t\pi_{1}+(1-t)\pi_{2}\middle|t\mathcal{M}(\pi_{1})+(1-t)\mathcal{M}(\pi_{2})\right)$
    $\qquad\leq t\textup{KL}\left(\pi_{1}\middle|\mathcal{M}(\pi_{1})\right)+(1-t)\textup{KL}\left(\pi_{2}\middle|\mathcal{M}(\pi_{2})\right),$
    which shows that $\mathcal{F}^{\eta}$ is convex in $\pi$ when $\eta=0$.
    For the case $\eta>0$ recall that
    $\mathcal{F}^{\eta}(\pi)$ $=\textup{KL}\left(\pi\middle|\varphi+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y+\eta\right),$
    and take $f(s)=s\log s$ with the convention $00\log 0=$ 0. Then, $\mathcal{F}^{\eta}(\pi)$ can be written as $\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\pi/\mathrm{d}\nu)\mathrm{d}\nu$ with $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x$ a $\sigma$-finite measure, while $\pi$ is a probability measure. Using Jensen’s inequality on the convex function $f$ w.r.t. the probability measure $\nu-\eta$ we have
    $\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\nu(\mathrm{d}x)$
    $\qquad=\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x$
    $\qquad\geq f\left(\int\limits_{\mathbb{R}^{d}}\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)\right)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x.$
    Since the integrand in the first term is always positive, the integral is in the support of $f$, and we can lower bound the first term with the minimum of $f$:
    $\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\nu(\mathrm{d}x)$ $\geq-e^{-1}+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x.$ (19)
    Now, we observe that
    $\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x$ (20)
    $=\int\limits_{\mathbb{R}^{d}}\!\!\frac{\eta\pi(x)}{\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta}\log\left(\frac{\pi(x)}{\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta}\right)\mathrm{d}x$
    $\geq\eta((1+\lambda)\mathtt{M}+\eta)^{-1}\mathcal{F}^{\eta}(\pi),$
    where we used Assumption 1 to obtain the lower bound. Using (19,20) and the definition of $\mathcal{F}^{\eta}$ we find
    $\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}+\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\mathcal{F}^{\eta}(\pi)$
    $\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}$
    $\mathcal{F}^{\eta}(\pi)\geq$ $-\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)^{-1}e^{-1}:=\mathtt{C}$
    The convexity of $\mathcal{F}^{\eta}$ follows from Lemma 1 with $\mu=\pi$ and $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x$ since the fact that $\eta>0$ guarantees that any probability measure with a Lebesgue density is absolutely continuous w.r.t. $\nu$.
  2. (b)
    To see that $\mathcal{F}_{\alpha}^{\eta}$ is proper observe that, since $\textup{KL}\left(\pi\middle|\pi_{0}\right)\geq 0$, $\mathcal{F}_{\alpha}^{\eta}(\pi)\geq\mathtt{C}$ for all $\pi\in\mathcal{P}(\mathbb{R}^{d})$.
    Now take a reference measure $\pi_{0}$ with finite entropy, i.e. $|\mathrm{H}(\pi_{0})|<+\infty$. We have that
    $\mathcal{F}_{\alpha}^{\eta}(\pi_{0})=-\mathrm{H}(\pi_{0})-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x.$
    In addition, under Assumption 1, for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}(\mathbb{R}^{d})$
    $|\log(\varphi(x)+\pi[\mathrm{k}(x,\cdot)]+\eta)|\leq\max(|\log(2\mathtt{M}+\eta)|,|\log(\eta)|),$
    showing that
    $-\int_{\mathbb{R}^{d}}\pi_{0}(x)\log(\varphi(x)+\pi_{0}[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}x\leq\max(|\log(2\mathtt{M}+\eta)|,|\log(\eta)|).$
    Hence, there exists $\pi=\pi_{0}$ for which $\mathcal{F}_{\alpha}^{\eta}(\pi)<+\infty$, showing that the functional is proper.
    Let $(\pi_{n})_{n\geq 1}\in(\mathcal{P}(\mathbb{R}^{d}))^{\mathbb{N}}$ be such that $\lim_{n\to+\infty}\pi_{n}=\pi\in\mathcal{P}(\mathbb{R}^{d})$. Since for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $|\mathrm{k}(x,y)|\leq\mathtt{M}$ and $\mathrm{k}$ is continuous, we have that $\lim_{n\to+\infty}\pi_{n}[\mathrm{k}(x,\cdot)]=\pi[\mathrm{k}(x,\cdot)]$ for each $x$. This and the fact that the Kullback–Leibler divergence is lower semi-continuous in both arguments [22, Lemma 1.4.3-(b)] guarantees that $\mathcal{F}^{\eta}$ is lower semi-continuous. [22, Lemma 1.4.3-(b)] also guarantees that $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is strictly convex and lower semi-continuous. It follows that $\mathcal{F}^{\eta}_{\alpha}$ is strictly convex and lower semi-continuous.
    To see that $\mathcal{F}^{\eta}_{\alpha}$ is coercive observe that $\mathcal{F}^{\eta}_{\alpha}$ is the sum of the lower bounded lower semi-continuous functional $\mathcal{F}^{\eta}$ and the coercive functional [22, Lemma 1.4.3-(c)] $\textup{KL}\left(\pi\middle|\pi_{0}\right)$, then for any $\beta\in\mathbb{R}$
    $S$ :=$:=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta\right\}$
    $\subseteq\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta-\mathtt{C}\right\}:=\tilde{S},$
    since $\mathcal{F}^{\eta}(\pi)\geq\mathtt{C}$. $\tilde{S}$ is relatively compact since $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is coercive and thus $S$ is also relatively compact, showing that $\mathcal{F}^{\eta}_{\alpha}$ is coercive.
Appendix B Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
We start by recalling the definition of subdifferentiability in Wasserstein spaces, see [4, Definition 10.1.1]. We denote by $\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ the space of probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ which are absolutely continuous w.r.t the Lebesgue measure and equip the space $\mathrm{L}^{2}(\mathbb{R}^{d},\pi):=\{f:\mathbb{R}^{d}\to\mathbb{R}^{d};\pi(\|f\|^{2})<\infty\}$ with the norm $\left\|f\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}:=\pi(\|f\|^{2})$.
Definition 1 (Fréchet subdifferential).
Let $\Phi:\ \mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ and $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, then $\xi\in\mathrm{L}^{2}(\mathbb{R}^{d},\pi)$ belongs to the strong Fréchet subdifferential $\partial_{\mathrm{s}}\Phi(\pi)$ of $\Phi$ at $\pi$ if for any sequence $(t_{n})_{n\geq 1}\in\mathrm{L}^{2}(\mathbb{R}^{d},\pi)$ such that $\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\to 0$ as $n\to\infty$ we have
$\liminf_{n\to\infty}\left.\left\{\Phi(t_{n\#}\pi)-\Phi(\pi)-\int\limits_{\mathbb{R}^{d}}\langle\xi(x),t_{n}(x)-x\rangle\mathrm{d}\pi(x)\right\}\middle/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\right.\geq 0.$
We are now ready to derive the subdifferential of $\mathcal{F}_{\alpha}^{\eta}$. First, let us denote by $\mathcal{G}^{\eta}(\pi):=-\int\limits_{\mathbb{R}^{d}}\log(\varphi(x)+\lambda\pi[\mathrm{k}(x,\cdot)]+\eta)\mathrm{d}\pi(x)$, so that $\mathcal{F}^{\eta}(\pi)=-\mathrm{H}(\pi)+\mathcal{G}^{\eta}(\pi)$.
Proposition 7 (Subdifferential of $\mathcal{G}^{\eta}$).
Under Assumption 1, the strong subdifferential of $\mathcal{G}^{\eta}$ is given by
$\partial_{\mathrm{s}}\mathcal{G}^{\eta}(\pi)=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right\}.$
Proof.
Let $\mathcal{L}:\ \mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{p}\to\left[0,+\infty\right)$ be given for any $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{d}$ by
$\mathcal{L}(\pi,x)=\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\varphi(x)$
and denote by $\xi_{1}(x,z):=\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)$ so that $\nabla\mathcal{L}(\pi,x)=\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z)$ for all fixed $\pi\in\mathcal{P}(\mathbb{R}^{d})$, where Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]) guarantees that we can swap the derivative w.r.t. $x$ with the integral w.r.t. $z$ since $\left\|\nabla_{1}\mathrm{k}\right\|\leq\mathtt{M}$ and $\mathrm{k}$ is bounded and thus integrable.
Further, define $\xi_{2}(x,z):=\lambda\nabla_{2}\mathrm{k}(x,z)$, let $g:\ \left[0,+\infty\right)\to\mathbb{R}$ be given for any $t\geq 0$ by $g(t)=-\log(t+\eta)$ and consider
$\mathcal{G}^{\eta}(t_{n\#}\pi)-\mathcal{G}^{\eta}(\pi)$ (21)
$-\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}\left[g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x)+g^{\prime}(\mathcal{L}(\pi,x))\xi_{1}(x,z)\right]\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)$
$=$ $\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(t_{n\#}\pi,x))\mathrm{d}t_{n\#}\pi(x)-\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}t_{n\#}\pi(x)$
$-\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x)\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)$
$+\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}t_{n\#}\pi(x)-\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}\pi(x)$
$-\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{1}(x,z)\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x).$
We consider two groups of terms separately.
We further decompose the first group of three terms in (21) into
$\int\limits_{\mathbb{R}^{d}}\!\!\left[g(\mathcal{L}(t_{n\#}\pi,x))-g(\mathcal{L}(\pi,x))-g^{\prime}(\mathcal{L}(\pi,x))\!\left(\int\limits_{\mathbb{R}^{d}}\!\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right)\!\right]\!\mathrm{d}t_{n\#}\pi(x)$ (22)
$+\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\mathrm{d}t_{n\#}\pi(x)$
$-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x)$
$\geq\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-y\rangle\mathrm{d}\pi(z)\right)\mathrm{d}t_{n\#}\pi(x)$
$\quad+\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\mathrm{d}t_{n\#}\pi(x)$
$\quad-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x)$
where the inequality follows since the convexity of $g$ implies
$g(\mathcal{L}(t_{n\#}\pi,x))-g(\mathcal{L}(\pi,x))-g^{\prime}(\mathcal{L}(\pi,x))\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)$
$\qquad\geq g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right).$
For the second group of terms in (21), we have, using Assumption 1 and the fact that $\mathcal{L}(\pi,x)$ is assumed to be positive for all $\pi\in\mathcal{P}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}$, we have, using the spectral norm induced by the Euclidean norm of vectors,
$\left\|\nabla^{2}g(\mathcal{L}(\pi,x))\right\|$ (23)
$\leq$ $\left\|\frac{(\mathcal{L}(\pi,x)+\eta)\left(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}^{2}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla^{2}\varphi(x)\right)}{(\mathcal{L}(\pi,x)+\eta)^{2}}\right\|$
$+\left\|\frac{(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))^{T}}{(\mathcal{L}(\pi,x)+\eta)^{2}}\right\|$
$\leq$ $\frac{((\lambda+1)\mathtt{M}+\eta)(\lambda+1)\mathtt{M}+(\lambda+1)^{2}\mathtt{M}^{2}}{\eta^{2}}:=\mathtt{C}_{0},$
where $\nabla\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)=\int\limits_{\mathbb{R}^{d}}\nabla^{2}_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)$ using Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]) since $\left\|\nabla_{1}^{2}\mathrm{k}\right\|\leq\mathtt{M}$ and $\xi_{1}$ is bounded and thus integrable. Using that $\left\|\nabla^{2}g(\mathcal{L}(\pi,x))\right\|\leq\mathtt{C}_{0}$ we get that for any $x_{1},x_{2}\in\mathbb{R}^{d}$,
$g(\mathcal{L}(\pi,x_{2}))\geq g(\mathcal{L}(\pi,x_{1}))+\langle g^{\prime}(\mathcal{L}(\pi,x_{1}))\int\limits_{\mathbb{R}^{d}}\xi_{1}(x_{1},z)\mathrm{d}\pi(z),x_{2}-x_{1}\rangle-\mathtt{C}_{0}\left\|x_{1}-x_{2}\right\|^{2},$
and [4, Proposition 10.4.2] gives that the limit as $n\to\infty$ of
$\frac{\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}(t_{n\#}\pi-\pi)(x)-\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,x))\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$ (24)
is non-negative. Combining (21)–(24) we obtain
$\mathcal{G}^{\eta}(t_{n\#}\pi)-\mathcal{G}^{\eta}(\pi)$ (25)
$-\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}\left[g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x)+g^{\prime}(\mathcal{L}(\pi,x))\xi_{1}(x,z)\right]\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)$
$\;\geq\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right)\mathrm{d}t_{n\#}\pi(x)$
$\;\phantom{\geq}+\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\mathrm{d}t_{n\#}\pi(x)$
$\;\phantom{\geq}-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x).$
Let us consider first
$\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\mathrm{d}t_{n\#}\pi(x)$
$\qquad-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x)$
$=$ $\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\mathrm{d}t_{n\#}\pi(x)$
$\qquad-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z),t_{n}(z)-z\right\rangle\mathrm{d}\pi(x)\mathrm{d}\pi(z)$
$=$ $\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x),t_{n}(z)-z\right\rangle\mathrm{d}\pi(z).$
Under Assumption 1, the function $x\mapsto g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)$ is continuous and bounded for any fixed $z$; in addition, $n\to\infty$ implies $t_{n\#}\pi\to\pi$ weakly, since $\mathbf{W}_{2}$ metrizes weak convergence. It follows that
$\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}t_{n\#}\pi(x)-\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}\pi(x)\to 0.$
The dominated convergence theorem then guarantees that
$\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x),t_{n}(z)-z\right\rangle\mathrm{d}\pi(z)\to 0$
as $n\to\infty$.
For the remaining term in (25), using that $\left\|\nabla^{2}\mathrm{k}\right\|\leq\mathtt{M}$ we get that for any $z_{1},z_{2}\in\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$
$\mathrm{k}(x,z_{2})\geq\mathrm{k}(x,z_{1})+\langle\nabla_{2}\mathrm{k}(x,z_{1}),z_{2}-z_{1}\rangle-\mathtt{M}\left\|z_{1}-z_{2}\right\|^{2}.$
Combing this and [4, Proposition 10.4.2] we get that for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the strong subdifferential of $\mathcal{L}(\pi,x)$ is given by $\xi_{2}(x,z)=\lambda\nabla_{2}\mathrm{k}(x,z)$, thus
$\lim_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$
In addition, since $\mathrm{k}$ is Lipschitz continuous with constant $\mathtt{M}$ we get that for any $\pi_{1},\pi_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{p}$
$|\mathcal{L}(\pi_{1},x)-\mathcal{L}(\pi_{2},x)|\leq\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}),$
and therefore
$\left|\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right|/\left\|t-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}$ (26)
$\qquad\qquad\qquad\leq\mathtt{M}+\int\limits_{\mathbb{R}^{d}}\left\|\xi_{2}(x,z)\right\|\mathrm{d}\pi(z)$
$\qquad\qquad\qquad\leq\mathtt{M}(1+\lambda).$
In particular, we have the following lower bound
$g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right)$
$\qquad\qquad\geq-\eta^{-1}\mathtt{M}(1+\lambda)\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}>-\infty.$
In addition, the function
$x\mapsto g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right)$
is Lipschitz continuous in $x$, as it is the product of Lipschitz continuous and bounded functions. In particular, $g^{\prime}(\mathcal{L}(\pi,x))$ is upper bounded by $\eta^{-1}$ and the Lipschitz continuity follows since its derivative is bounded (23). Similarly, the function
$x\mapsto\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)$
is upper bounded by (26) and Lipschitz continuous (this follows from the defintion of $\mathcal{L}$ and Assumption 1). Hence, the sequence of functions is equicontinuous and we can apply Fatou’s Lemma for weakly converging measures [24, Theorem 4.1] to obtain
$\liminf_{n\to\infty}\frac{\int\limits_{\mathbb{R}^{d}}\!g^{\prime}(\mathcal{L}(\pi,x))\!\left(\!\mathcal{L}(t_{n\#}\pi,x)\!-\!\mathcal{L}(\pi,x)\!-\!\int\limits_{\mathbb{R}^{d}}\!\langle\xi_{2}(x,z),t_{n}(z)\!-\!z\rangle\mathrm{d}\pi(z)\!\right)\!\mathrm{d}t_{n\#}\pi(x)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}\geq 0,$
which gives the result. ∎
[4, Theorem 10.4.9] shows that the subdifferential of $\mathcal{M}(\pi)=\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is given by
$\partial_{\mathrm{s}}\mathcal{M}(\pi)=\{x\mapsto\nabla\log(\mathrm{d}\pi/\mathrm{d}\pi_{0})(x)\}.$ (27)
Similarly, the subdifferential of $\mathrm{H}(\pi):=-\int\limits_{\mathbb{R}^{d}}\log(\pi(x))\mathrm{d}\pi(x)$ is
$\partial_{\mathrm{s}}\mathrm{H}(\pi)=\{x\mapsto\nabla\log\pi(x)\}.$ (28)
Combining the above with Proposition 7 we obtain the subdifferential for $\mathcal{F}_{\alpha}^{\eta}$
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ =$=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right.$
$\left.\qquad\qquad+\nabla\log\pi(x)+\alpha\nabla\log(\mathrm{d}\pi/\mathrm{d}\pi_{0})(x)\right\}.$
We can then apply the definition of a Wasserstein gradient flow [4, Definition 11.1.1] to establish that if there exists $(\pi_{t})_{t\geq 0}$ such that for any $t\in(0,+\infty)$, $\pi_{t}$ admits a density w.r.t. the Lebesgue measure and (8) holds in the sense of distributions then $(\pi_{t})_{t\geq 0}$ is a Wasserstein gradient flow associated with $\mathcal{F}_{\alpha}^{\eta}$.
Appendix C Proof on MKVSDE
C.1 Proof of Proposition 2
To prove existence and uniqueness of the solution of (9) we use standard tools for McKean–Vlasov processes. In particular, we only need to show that the drift of (9), denoted by $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, given for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ by
$b(x,\pi):=\int\limits_{\mathbb{R}^{d}}b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x),$ (29)
is Lipschitz continuous. The remainder of the proof is classical and is omitted, see for instance [51, Theorem 1.1] or [16, Appendix B.2] for a recent result on a similar scheme.
Let $\alpha,\eta>0$. First, we show that $b^{\eta}$ in (10) is Lipschitz continuous. Under Assumption 1, the forcing term $\varphi$, $\mathrm{k}$ and their gradients are Lipschitz continuous, hence, we have for any $x_{1},x_{2}\in\mathbb{R}^{d}$, $z\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}_{1}(\mathbb{R}^{d})$
$\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|$
$\leq$ $\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z,x_{1})}{\lambda\pi_{1}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z,x_{2})}{\lambda\pi_{2}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x_{1},z)+\nabla\varphi(x_{1})}{\lambda\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]+\varphi(x_{1})+\eta}-\frac{\lambda\nabla_{1}\mathrm{k}(x_{2},z)+\nabla\varphi(x_{2})}{\lambda\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]+\varphi(x_{2})+\eta}\right\|$
$\leq$ $\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z,x_{1})-\lambda\nabla_{2}\mathrm{k}(z,x_{2})}{\lambda\pi_{1}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z,x_{2})}{\lambda\pi_{1}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z,x_{2})}{\lambda\pi_{2}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x_{1},z)+\nabla\varphi(x_{1})-\lambda\nabla_{1}\mathrm{k}(x_{2},z)-\nabla\varphi(x_{2})}{\lambda\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]+\varphi(x_{1})+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x_{2},z)+\nabla\varphi(x_{2})}{\lambda\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]+\varphi(x_{1})+\eta}-\frac{\lambda\nabla_{1}\mathrm{k}(x_{2},z)+\nabla\varphi(x_{2})}{\lambda\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]+\varphi(x_{2})+\eta}\right\|$
$\leq$ $\lambda(\mathtt{M}/\eta)\left\|x_{1}-x_{2}\right\|+\lambda^{2}(\mathtt{M}/\eta^{2})|\pi_{2}\left[\mathrm{k}(z,\cdot)\right]-\pi_{1}\left[\mathrm{k}(z,\cdot)\right]|$
$+(\lambda+1)(\mathtt{M}/\eta)\left\|x_{1}-x_{2}\right\|$
$+(\lambda+1)(\mathtt{M}/\eta^{2})(\mathtt{M}\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|).$
Hence, we have
$\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|),$ (30)
with $\mathtt{C}_{1}$ depending upon $\lambda,\eta$ and $\mathtt{M}$.
Following a similar approach, we can show that $b^{\eta}$ is Lipschitz continuous in $z$: for any $z_{1},z_{2}\in\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$ and $\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$
$\left\|b^{\eta}(x,z_{1},\nu)-b^{\eta}(x,z_{2},\nu)\right\|$ $\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{1})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}-\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{2})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right\|$
$\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)-\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$+\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$+\lambda(\mathtt{M}/\eta^{2})|\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})-\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]-\varphi(z_{1})|$
$\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|+\lambda(\mathtt{M}^{2}/\eta^{2})(1+\lambda)\left\|z_{1}-z_{2}\right\|$
$\leq\mathtt{C}_{2}\left\|z_{1}-z_{2}\right\|,$
for $\mathtt{C}_{2}$ which depends on $\lambda,\eta$ and $\mathtt{M}$.
Using (30), for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ we have
$\left\|\int b^{\eta}(x_{1},z,\pi_{1})\mathrm{d}\pi_{1}(z)-\int b^{\eta}(x_{2},z,\pi_{2})\mathrm{d}\pi_{2}(z)\right\|$ (31)
$\qquad\leq\int\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|\mathrm{d}\pi_{1}(z)$
$\qquad+|\int b^{\eta}(x_{2},z,\pi_{2})\mathrm{d}(\pi_{1}-\pi_{2})(z)|$
$\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)$
$\qquad+|\pi_{2}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]-\pi_{1}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]|.$
Using the dual representation of $\mathbf{W}_{1}$ and the Lipschitz continuity of $\mathrm{k}$ we also have that for any $x_{1},x_{2}\in\mathbb{R}^{d}$
$|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|$ $\leq|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{2},\cdot)\right]|$
$+|\pi_{1}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|$
$\leq\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\mathtt{M}\left\|x_{1}-x_{2}\right\|,$
and
$|\pi_{2}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]-\pi_{1}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]|\leq\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$ (32)
Using the above, for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ we have
$\left\|\int b^{\eta}(x_{1},z,\pi_{1})\mathrm{d}\pi_{1}(z)-\int b^{\eta}(x_{2},z,\pi_{2})\mathrm{d}\pi_{2}(z)\right\|$ (33)
$\qquad\leq\mathtt{C}_{1}((1+\mathtt{M})\left\|x_{1}-x_{2}\right\|+\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}))+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$
Using Assumption 2, it then follows that
$\left\|b(x_{1},\pi_{1})-b(x_{2},\pi_{2})\right\|$ $\leq\mathtt{C}_{1}((1+\mathtt{M})\left\|x_{1}-x_{2}\right\|+\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}))$ (34)
$\phantom{\leq}\quad+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
$\leq(\mathtt{C}_{1}+\mathtt{C}_{1}\mathtt{M}+\mathtt{C}_{2}+\alpha\mathtt{L})(\left\|x_{1}-x_{2}\right\|+\mathbf{W}_{1}(\pi_{1},\pi_{2})).$
C.2 Proof of Proposition 3
Under Assumption 1, Proposition 1(a) ensures that the functional $\mathcal{F}^{\eta}$ is convex and lower bounded. In addition, Assumption 2 guarantees that $U$ is sufficiently regular to satisfy [30, Assumption 2.2]. To see this, we can use Assumption 2(c) and use the Cauchy-Schwarz inequality to obtain
$\langle\nabla U(x),x\rangle$ $=\langle\nabla U(x)-\nabla U(0),x\rangle+\langle\nabla U(0),x\rangle\geq\mathtt{m}\left\|x\right\|^{2}-\mathtt{c}-\left\|x\right\|\left\|\nabla U(0)\right\|.$
Then, we use the following version of Young’s inequality, $\alpha\beta\leq\alpha^{2}/(2\epsilon)+\beta^{2}\epsilon/2$ for all $\epsilon>0$, to show that
$-\left\|x\right\|\left\|\nabla U(0)\right\|\geq-\left\|x\right\|^{2}/(2\epsilon)-\epsilon\left\|\nabla U(0)\right\|^{2}/2.$
It follows that for all $\epsilon>1/(2\mathtt{m})$ we have, for any $x\in\mathbb{R}^{d}$, $\langle\nabla U(x),x\rangle\geq\mathtt{a}\left\|x\right\|^{2}+\mathtt{b}$, where $\mathtt{a}:=\mathtt{m}-1/(2\epsilon)>0,\mathtt{b}=-\mathtt{c}-\epsilon\left\|\nabla U(0)\right\|^{2}/2$.
As shown in (33), the first component of the drift of the MKVSDE (9) is Lipschitz continuous. In addition, using the fact that under Assumption 1 $\mathrm{k},\varphi\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]), we have that $b(x,\nu)$ in (29) is $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ for all fixed $\nu\in\mathcal{P}(\mathbb{R}^{d})$. Finally, we need to show that
$\nabla b(x,\nu)$ $:=\int\limits_{\mathbb{R}^{d}}\nabla_{1}b^{\eta}(x,z,\nu)\mathrm{d}\nu(z)=\nabla b^{1}(x,\nu)+\nabla b^{2}(x,\nu),$
where we defined
$\nabla b^{1}(x,\nu):=$ $\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x)}{\lambda\nu\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z),$
and
$\nabla b^{2}(x,\nu):=$ $\int\limits_{\mathbb{R}^{d}}\frac{(\lambda\nabla_{1}^{2}\mathrm{k}(x,z)+\nabla^{2}\varphi(x))(\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta)}{(\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta)^{2}}\mathrm{d}\nu(z)$
$-\int\limits_{\mathbb{R}^{d}}\frac{(\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x))(\lambda\nu\left[\nabla_{1}\mathrm{k}(x,\cdot)\right]+\nabla\varphi(x))}{(\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta)^{2}}\mathrm{d}\nu(z),$
is jointly continuous in $(x,\nu)$. To do so, consider a sequence $(x_{n},\nu_{n})_{n\geq 0}\in(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d}))^{\mathbb{N}}$ such that such that $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)\in\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$. Then,
$\left\|\nabla b^{1}(x_{n},\nu_{n})-\nabla b^{1}(x,\nu)\right\|$
$\leq$ $\left\|\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu_{n}(z)-\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)\right\|$
$+\left\|\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)-\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x)}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)\right\|$
$+\left\|\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x)}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)-\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x)}{\lambda\nu\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)\right\|$
$\leq$ $\left\|\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu_{n}(z)-\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)\right\|$
$+\frac{\lambda}{\eta}\left\|x-x_{n}\right\|+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\mathbf{W}_{1}(\nu_{n},\nu),$
where we used the fact that $\mathrm{k}$ is Lipschitz continuous with Lipschitz constant $\mathtt{M}$ and, for the last term, the dual representation of the $\mathbf{W}_{1}$ distance. In addition, since $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and $\left\|\mathrm{d}^{3}\mathrm{k}(x,y)\right\|\leq\mathtt{M}$, the function
$z\mapsto\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}$
is Lipschitz continuous with Lipschitz constant $\mathtt{M}_{1}$ for all $n$ uniformly in $x_{n}$. Hence, using again the dual representation of the $\mathbf{W}_{1}$ distance, we find
$\left\|\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu_{n}(z)-\int\limits_{\mathbb{R}^{d}}\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x)}{\lambda\nu\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\mathrm{d}\nu(z)\right\|$
$\qquad\leq\frac{\lambda}{\eta}\left\|x-x_{n}\right\|+\left(\mathtt{M}_{1}+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\right)\mathbf{W}_{1}(\nu_{n},\nu).$
Similarly, we can show that the second term satisfies
$\left\|\nabla b^{2}(x_{n},\nu_{n})-\nabla b^{2}(x,\nu)\right\|\leq C(\left\|x-x_{n}\right\|+\mathbf{W}_{1}(\nu_{n},\nu))$
for some finite $C$.
Since $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)$ implies $\left\|x-x_{n}\right\|\to 0$ and $\mathbf{W}_{1}(\nu_{n},\nu)\to 0$, we have that $\nabla b(x,\nu)$ is jointly continuous. Then, the result follows directly from [30, Theorem 2.11].
Appendix D Particle System and Time Discretization
D.1 Proof of Proposition 4
For any $\{x_{1}^{k,N}\}_{k=1}^{N},\{x_{2}^{k,N}\}_{k=1}^{N}\in(\mathbb{R}^{d})^{N}$ and $\ell\in\{1,\dots,N\}$, let us denote $\pi_{1}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{1}^{k,N}}$ and $\pi_{2}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{2}^{k,N}}$. Using (34) and the fact that for empirical measures
$\mathbf{W}_{1}(\pi_{1}^{N},\pi_{2}^{N})\leq N^{-1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|$
we have that
$\left\|b(x_{1}^{\ell,N},\pi_{1}^{N})-b(x_{2}^{\ell,N},\pi_{2}^{N})\right\|$ $\leq 2(\mathtt{C}_{1}(1+\mathtt{M})+\mathtt{C}_{2}+\alpha\mathtt{L})\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$
The existence and strong uniqueness of a solution to (11) is then a straightforward consequence of the above and [36, Chapter 5, Theorem 2.9 and 2.5].
The propagation of chaos result is a straightforward consequence of the Lipschitz continuity of $b$. Combining (31), (32) and Assumption 2 we have for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}_{1}(\mathbb{R}^{d})$
$\left\|b(x_{1},\pi_{1})-b(x_{2},\pi_{2})\right\|$ (35)
$\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
$\qquad\leq(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2})|).$
The rest of the proof is classical [51] and given for completeness, see [16, Appendix B.3] for a proof in a similar scenario.
Using (35), we have for any $t\geq 0$
$\mathbb{E}\left[\sup_{s\in[0,t]}\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right]$ (36)
$\qquad\leq\int\limits_{0}^{t}\mathbb{E}\left[\left\|b(\bm{X}_{s},\pi_{s})-b(\bm{X}_{s}^{1,N},\pi^{N}_{s})\right\|\right]\mathrm{d}s$
$\qquad\leq(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]\mathrm{d}s$
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|\right]\mathrm{d}s$
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})\mathrm{d}s$
$\qquad\leq 2(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]\mathrm{d}s$
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|\right]\mathrm{d}s,$
where the last inequality follows from the fact that
$\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})\leq\mathbf{W}_{2}(\pi_{s},\pi_{s}^{N})\leq\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|.$ (37)
Now, consider $N$ independent copies of the nonlinear process $\bm{X}_{s}$, $\{(\bm{X}_{s}^{k,\star})_{t\geq 0}\}_{k=1}^{N}$. We can bound the second term in the above with
$\mathbb{E}\left[\left|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]$
$\qquad\leq\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})\right|\right]$
$\qquad+\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]$
$\qquad\leq\mathtt{M}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]$
$\qquad+\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]$
where we used the Lipschitz continuity of $\mathrm{k}$, (37) and the fact that $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is exchangeable to obtain the last inequality. Plugging the above into (36) and using Jensen’s inequality we obtain
$\mathbb{E}\left[\sup_{s\in[0,t]}\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right]$ $\leq(2+\mathtt{M})(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]\mathrm{d}s$
$+\frac{\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L}}{N}\int_{0}^{t}\mathbb{E}\left[\sum_{i=1}^{N}|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|^{2}\right]^{1/2}\mathrm{d}s.$
Using Popoviciu’s inequality on variances [47] and recalling that $0\leq\mathrm{k}(x,y)\leq\mathtt{M}$ for all $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, we have
$\mathbb{E}\left[|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|^{2}\mid\bm{X}_{s}^{1,N}\right]$ $=\textrm{var}\left(\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})\mid\bm{X}_{s}^{1,N}\right)$
$\leq\mathtt{M}^{2}/4.$
It follows that
$\mathbb{E}\left[\sup_{s\in[0,t]}\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right]$ $\leq(2+\mathtt{M})(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]\mathrm{d}s$
$+\mathtt{M}/2(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})N^{-1/2}t.$
Using Grönwall’s lemma we get that for any $T\geq 0$ there exists $C_{N}(T)\geq 0$ such that for any $N\in\mathbb{N}$
$\mathbb{E}\left[\sup_{t\in[0,T]}\left\|\bm{X}_{t}-\bm{X}_{t}^{1,N}\right\|\right]\leq C_{N}(T)N^{-1/2},$
which concludes the proof.
D.2 Proof of Proposition 5
First, we recall that $b$ is given for any $x\in\mathbb{R}^{d}$, $\nu\in\mathcal{P}(\mathbb{R}^{d})$ and $y\in\mathbb{R}^{p}$ by
$b(x,\nu)=\int b^{\eta}(\bm{X}_{t},z,\pi_{t})\mathrm{d}\pi_{t}(z)-\alpha\nabla U(\bm{X}_{t})$
with $b^{\eta}$ defined in (10). Recall that under Assumption 1 and 2 we have (34), i.e., there exist $C_{1}\geq 0$ such that for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$
$\left\|b(x_{1},\pi_{1})-b(x_{2},\pi_{2})\right\|\leq C_{1}\left(\left\|x_{1}-x_{2}\right\|+\mathbf{W}_{1}(\pi_{1},\pi_{2})\right).$ (38)
Furthermore, using Assumption 1 we have for any $x_{1},x_{2}\in\mathbb{R}^{d}$, $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$ and $y\in\mathbb{R}^{p}$,
$\left\|b^{\eta}(x,z,\nu)\right\|$ $\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\nu\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}\right\|+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right\|$ (39)
$\leq\lambda\mathtt{M}/\eta+\lambda(\mathtt{M}+1)/\eta:=C_{0}.$
In addition, note that for any $x_{1}^{1:N},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ we have for any $i\in\{1,2\}$
$\mathbf{W}_{1}(\pi_{1}^{N},\pi_{2}^{N})\leq\frac{1}{N}\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|,\qquad\pi_{i}^{N}=\frac{1}{N}\sum_{k=1}^{N}\delta_{x^{k,N}}.$ (40)
Let $N\in\mathbb{N}$ and denote $B_{N}:(\mathbb{R}^{d})^{N}\to(\mathbb{R}^{d})^{N}$ given for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$ by
$B_{N}(x^{1:N})=\{b(x^{k,N},\pi^{N})\}_{k\in\{1,\dots,N\}},\qquad\pi^{N}=\frac{1}{N}\sum_{k=1}^{N}\delta_{x^{k,N}}.$
Therefore, using (38)–(40) we have for any $x^{1:N}_{1},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$
$\left\|B_{N}(x^{1:N}_{1})-B_{N}(x^{1:N}_{2})\right\|$ (41)
$\qquad\leq C_{1}\left(\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|^{2}+N\mathbf{W}_{1}\left(\frac{1}{N}\sum_{k=1}^{N}\delta_{x_{1}^{k,N}},\frac{1}{N}\sum_{k=1}^{N}\delta_{x_{2}^{k,N}}\right)^{2}\right.$
$\qquad\qquad\left.+2\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|\mathbf{W}_{1}\left(\frac{1}{N}\sum_{k=1}^{N}\delta_{x_{1}^{k,N}},\frac{1}{N}\sum_{k=1}^{N}\delta_{x_{2}^{k,N}}\right)\right)^{1/2}$
$\qquad\leq C_{1}\left(\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|^{2}+\frac{1}{N}\left(\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|\right)^{2}\right.$
$\qquad\qquad\left.+2/N\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|\right)^{1/2}$
$\qquad\leq 2C_{1}\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|$
$\qquad\leq 2C_{1}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$ (42)
Using (38)–(39) and Assumption 2(c) we have for any $x^{1:N}_{1},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$
$\langle B_{N}(x^{1:N}_{1})-B_{N}(x^{1:N}_{2}),x_{1}^{1:N}-x_{2}^{1:N}\rangle$ (43)
$\qquad\leq-\alpha\mathtt{m}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}+\alpha\mathtt{c}N+2C_{0}\sum_{k=1}^{N}\left\|x_{1}^{k,N}-x_{2}^{k,N}\right\|$
$\qquad\leq-\alpha\mathtt{m}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}+\alpha\mathtt{c}N+2C_{0}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$ (44)
Let $R=\max(4C_{0}N^{1/2}/(\alpha\mathtt{m}),(2\mathtt{c}N/\mathtt{m})^{1/2})$ with $\mathtt{m}$ and $\mathtt{c}$ as in Assumption 2(c). Then, for any $x^{1:N}_{1},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ with $\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|\geq R$ we further have
$\langle B_{N}(x^{1:N}_{1})-B_{N}(x^{1:N}_{2}),x_{1}^{1:N}-x_{2}^{1:N}\rangle$ $\leq-(\alpha\mathtt{m}/2)\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}.$ (45)
We conclude upon combining (41), (43) and [19, Corollary 2].
D.3 Proof of Proposition 6
Let $\alpha,\eta>0$. First we show that $\{\pi^{N}\}_{N\in\mathbb{N}}$ is relatively compact in $\mathcal{P}_{1}(\mathbb{R}^{d})$. Let $N\in\mathbb{N}$ and assume that $\bm{X}_{0}^{1:N}=0$. Define for any $t\geq 0$
$\bm{M}_{t}^{N}:=(1/2)\left\|\bm{X}_{t}^{1,N}\right\|^{2}-\int\limits_{0}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+\alpha d\right\}\mathrm{d}u,$
where $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is given in (11). Using Itô’s formula we have that
$(1/2)\left\|\bm{X}_{t}^{1,N}\right\|^{2}=(1/2)\left\|\bm{X}_{s}^{1,N}\right\|^{2}+\int\limits_{s}^{t}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u+d\alpha(t-s).$
Denote by $(\mathcal{F}_{t}^{N})_{t\geq 0}$ the natural filtration of $\{(\bm{B}_{t}^{k})_{t\geq 0}\}_{k=1}^{N}$. Then, $\mathbb{E}\left[\bm{M}_{t}^{N}\mid\mathcal{F}_{s}^{N}\right]=\bm{M}_{s}^{N}$, showing that $\bm{M}_{t}^{N}$ is an $\mathcal{F}^{N}$-local martingale.
Since $\mathbb{E}\left[\bm{M}_{t}^{N}-\bm{M}_{s}^{N}\mid\mathcal{F}_{s}^{N}\right]=0$, we also have that
$(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-(1/2)\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+\alpha d\right\}\mathrm{d}u\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},\int\limits_{\mathbb{R}^{d}}b^{\eta}(\bm{X}_{u}^{1,N},z,\pi^{N}_{u})\mathrm{d}\pi_{u}^{N}(z)\rangle\right\}\mathrm{d}u\right]$
$\qquad-\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},\alpha\nabla U(\bm{X}_{u}^{1,N})\rangle-\alpha d\right\}\mathrm{d}u\right],$
where we used the definition of $b$ in (29).
Using Assumption 2(c) with $x_{1}=x$ and $x_{2}=0$, we have for any $x\in\mathbb{R}^{d}$
$\langle\nabla U(x),x\rangle\geq\langle\nabla U(0),x\rangle+\mathtt{m}\left\|x\right\|^{2}-\mathtt{c}.$
Therefore, using this result and the Cauchy–Schwarz inequality we obtain that for any $t\geq 0$
$(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-(1/2)\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
$\qquad\leq\mathbb{E}\left[\int\limits_{s}^{t}\left\{\int\limits_{\mathbb{R}^{d}}\left\|b^{\eta}(\bm{X}_{u}^{1,N},z,\pi^{N}_{u})\right\|\left\|\bm{X}_{u}^{1,N}\right\|\mathrm{d}\pi_{u}^{N}(z)\right\}\mathrm{d}u\right]$
$\qquad+\mathbb{E}\left[\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\left\|\bm{X}_{u}^{1,N}\right\|^{2}+\alpha\left\|\nabla U(0)\right\|\left\|\bm{X}_{u}^{1,N}\right\|+\alpha\mathtt{c}+\alpha d\right\}\mathrm{d}u\right]$
$\qquad=\int\limits_{s}^{t}\mathbb{E}\left[\int\limits_{\mathbb{R}^{d}}\left\|b^{\eta}(\bm{X}_{u}^{1,N},z,\pi^{N}_{u})\right\|\left\|\bm{X}_{u}^{1,N}\right\|\mathrm{d}\pi_{u}^{N}(z)\right]\mathrm{d}u$
$\qquad+\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|\right]+\alpha\mathtt{c}+\alpha d\right\}\mathrm{d}u,$
where the last line follows from Tonelli’s Theorem since all integrated functions are positive (or always negative, in which case we can consider minus the integral itself).
Let $\mathcal{V}_{t}^{N}=(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]$. Appealing to the fundamental theorem of calculus we get that
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\mathbb{E}\left[\int\limits_{\mathbb{R}^{d}}\left\|b^{\eta}(\bm{X}_{t}^{1,N},z,\pi^{N}_{t})\right\|\left\|\bm{X}_{t}^{1,N}\right\|\mathrm{d}\pi_{t}^{N}(z)\right]$
$-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]+\alpha\mathtt{c}+\alpha d$
Using Jensen’s inequality we have $\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]\leq\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]^{1/2}$ and using that, as shown in (39), for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}(\mathbb{R}^{d})$ we have $\left\|b^{\eta}(x,z,\pi)\right\|\leq C_{0}$ we have:
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $C_{0}(\mathcal{V}_{t}^{N})^{1/2}-\alpha\mathtt{m}\mathcal{V}_{t}^{N}+\alpha\left\|\nabla U(0)\right\|(\mathcal{V}_{t}^{N})^{1/2}+\alpha\mathtt{c}+\alpha d$
$=$ $(C_{0}+\alpha\left\|\nabla U(0)\right\|)(\mathcal{V}_{t}^{N})^{1/2}+\alpha\mathtt{c}+\alpha d-\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
Noting that for any $\mathsf{a},\mathsf{b}$ such that $\mathsf{ab}\geq 1/2$, for any $x\geq 0$ we have $\sqrt{x}\leq\mathsf{a}+\mathsf{b}x$, and setting $\mathsf{a}=(C_{0}+\alpha\left\|\nabla U(0)\right\|)/\alpha\mathtt{m}$ and $\mathsf{b}=1/2\mathsf{a}$ we have:
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)^{2}}{\alpha\mathtt{m}}+\alpha\mathtt{c}+\alpha d-\frac{1}{2}\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
Hence, for any $t\geq 0$ and any $N\in\mathbb{N}$ we get that $\mathcal{V}_{t}^{N}\leq C$ with
$C=2\left[\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)^{2}}{\alpha\mathtt{m}}+\alpha\mathtt{c}+\alpha d\right]/\alpha\mathtt{m}.$ (46)
Therefore, letting $t\to+\infty$ we get that for any $N\in\mathbb{N}$, $\int\limits_{\mathbb{R}^{d}}\left\|x\right\|^{2}\mathrm{d}\pi^{N}(x)\leq C$. Hence, $\{\pi^{N}\}_{N\in\mathbb{N}}$ is relatively compact in $\mathcal{P}_{1}(\mathbb{R}^{d})$ using [4, Proposition 7.1.5].
Let $\pi^{\star}$ be a cluster point of $\{\pi^{N}\}_{N\in\mathbb{N}}$. Le us denote by $\pi_{t}(\pi^{\star}),\pi_{t}(\pi^{\star})^{N}$ the law of $\bm{X}_{t}$ following (9) with initial condition $\bm{X}_{0}\sim\pi^{\star}$ and $\bm{X}_{t}^{1,N}$ following (11) with initial condition $\bm{X}_{0}^{1,N}\sim\pi^{\star}$. Let $(N_{k})_{k\in\mathbb{N}}$ be an increasing sequence such that $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. We have that for any $t\geq 0$
$\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})$ $\leq\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k}})$ (47)
$+\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))+\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))$
$+\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))+\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star}).$
We now control each of these terms. Let $\varepsilon>0$, for sufficiently large $t\geq 0$ Proposition 3 ensures $\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star})\leq\varepsilon$. Using the Lipschitz continuity of the drift established in (31) and Assumption 2 with [16, Lemma 20], there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and any $t\in[0,T]$ we have $\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))\leq C_{T}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Using Proposition 4, there exists $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$ we have that $\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))\leq\varepsilon$. Since $\pi^{N_{k}}$ is invariant for $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ we get that $\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))=0$. Finally, there exists $k_{2}\in\mathbb{N}$ such that for any $k\geq k_{2}$ we have $\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k_{2}}})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Combining these results in (47), we get that $\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})\leq 4\varepsilon$. Therefore, since $\varepsilon>0$ is arbitrary, we have that $\pi^{\star}=\pi_{\alpha,\eta}^{\star}$, which concludes the proof.